OpenStudy (anonymous):

Could someone show me how to do: cd^3/c^3d^9?

OpenStudy (anonymous):

\[\frac{ cd^3 }{ c^3d ^{-9}}\]

OpenStudy (akashdeepdeb):

When like terms [here, c& d] are divided, their powers are subtracted. \[\frac{ cd^3 }{ c^3d^{-9} } = c^{1-3}.d^{3-(-9)} = c^{-2}.d^{12}\] Understood? :)

OpenStudy (anonymous):

Thank you!

OpenStudy (akashdeepdeb):

:)